OpenStudy (kj4uts):

The profit function for a firm making DVDs is P(x) = 88x - x^2 -1200. Find the number of DVDs at which maximum profit is achieved, and find the maximum profit. Please explain. Thank you!

OpenStudy (kj4uts):

@peachpi

OpenStudy (some.random.cool.kid):

@Nnesha

OpenStudy (anonymous):

Is x the number of dvds?

OpenStudy (kj4uts):

Yeah I think so.

OpenStudy (anonymous):

It looks like this is asking for the vertex of the parabola. \(x=-\frac{b}{2a}\) is the amount of dvd where profit is max \(P(-\frac{b}{2a})\) is the maximum profit

OpenStudy (kj4uts):

Do I plug in the 88 and -1200 in that formula?

OpenStudy (anonymous):

-b/(2a) = -88/(2*-1)

OpenStudy (anonymous):

a is the coefficient of x² b is the coefficient of x c is the constant It helps to write it in descending order

OpenStudy (kj4uts):

44 so that would be the max number of dvds 88(44)-44^2-1200=$736 (the max profit)

OpenStudy (kj4uts):

@peachpi thank you :)

OpenStudy (anonymous):

that's what I got :) you're welcome